87,102,105,111,114 . find the next number?
Answer by jim_thompson5910(35256)

Notice how the digits of 87 add to 8+7 = 15

If you add 15 to 87, you get 87+15 = 102, which is the next term (how curious...)

-------------------------------------------------------

If you add the digits of 102, you get 1+0+2 = 3

Add this to 102 to get 102+3 = 105, which is the next term (possibly a coincidence, we should check all the terms given to find out)

See the pattern?


-------------------------------------------------------

If you add the digits of 105, you get 1+0+5 = 6

Add this to 102 to get 105+6 = 111, which is again the next term.

-------------------------------------------------------

If you add the digits of 111, you get 1+1+1 = 3

Add this to 102 to get 111+3 = 114, which is the next term

-------------------------------------------------------
So if we assume that the pattern keeps up, then...

The digits of 114 add to 1+1+4 = 6

Add this to 114 to get 114+6 = 120

=======================================================

Answer:

So the next number is 120